### v2

@@ 3 Posterior smoothing with dimension-independent error
Since MLwG requires the gradient of the log posterior density in Eqs. 11 and 12, we propose to approximate the non-smooth posterior density $\pi$ in Eq. 8 by a smooth density $\pi_{\varepsilon}$. Moreover, we show that the introduced error between $\pi$ and $\pi_{\varepsilon}$ is distributed uniformly, leading to a local dimension-independent error on the marginal distribution of any block $\mathbf{x}_{i}$.
The non-smoothness of the posterior Eq. 8 originates from the potential $\varphi_{0}(\mathbf{x})={\rm TV}(\mathbf{x})$ in Eq. 4. Hence, we replace $\varphi_{0}$ with a smoothed potential $\varphi_{\varepsilon}$ for some small $\varepsilon>0$, such that $\varphi_{\varepsilon}\to\varphi_{0}$ as $\varepsilon\to 0$. Various smoothing methods exist, but it is crucial to ensure that the introduced error remains small, particularly when $\mathbf{x}$ resides in a high-dimensional space. In this paper, we consider the approximation
(13) $\varphi_{\varepsilon}(\mathbf{x}):=\delta\sum_{\alpha=1}^{d}\sqrt{(\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon},$
@@
(14) $\pi_{\varepsilon}(\mathbf{x})\propto\exp\left(-\frac{\lambda}{2}\|\mathbf{y}-\mathbf{A}\mathbf{x}\|_{2}^{2}-\delta\sum_{\alpha=1}^{d}\sqrt{(\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon}\right).$
As we modify $\varphi_{0}$, a function in $\mathbb{R}^{d}$, the distance between $\pi$ and $\pi_{\varepsilon}$ generally depends on the dimension $d$, for instance, one can show $\text{KL}(\pi|\pi_{\varepsilon})=\mathcal{O}(d\varepsilon)$. Here $\text{KL}(\pi|\pi_{\varepsilon})$ is the Kullback-Leibler divergence of $\pi$ from $\pi_{\varepsilon}$, see, e.g., [25]. However, when examining the marginals of $\pi$ and $\pi_{\varepsilon}$ over small blocks $\mathbf{x}_{i}$, we can show that the approximation error is independent of the full dimension, and thus dimension-independent. We comment that such dimension independence is crucial for solving the image deblurring problem. It ensures that the smoothing error is evenly distributed across the image, rather than concentrating on certain pixels and creating unwanted artifacts in the image. Ensuring a uniformly distributed error is not a concept unique to image reconstruction, for other applications see [7, 33, 12].
The key to achieve such results lies in the sparse conditional structure. Intuitively, such local structure makes the distribution on small blocks mostly influenced by local modifications to $\varphi_{0}$, and modifications on remote blocks have little impact on it. To rigorously justify the above observations, we introduce the concept of $c$-diagonal block dominance, which is imposed on $\mathbf{C}=\mathbf{A}^{\rm T}\mathbf{A}$, to quantify how the blurring matrix $\mathbf{A}$ is locally concentrated.
@@
Consider a positive definite matrix $\mathbf{C}\in\mathbb{R}^{d\times d}$ and let $\mathbf{C}_{ij}$ denote the $(i,j)$-th sub block of $\mathbf{C}$ of size $q$. $\mathbf{C}$ is called $c$-diagonal block dominant for some $c>0$, if there exists a symmetric matrix $M\in\mathbb{R}^{b\times b}$ s.t.
@@
[32] introduces a similar blockwise log-concavity condition, which is crucial for the dimension-independent convergence rate therein. The $c$-diagonal block dominance here can be viewed as a $\ell_{1}$ version of the blockwise log-concavity condition (which is a $\ell_{2}$ condition).
Now we state the main theorem that shows that the marginal error of $\pi$ and $\pi_{\varepsilon}$ is dimension-independent if $\mathbf{A}^{\rm T}\mathbf{A}$ is $c$-diagonal block dominant.
@@
Consider the target distribution $\pi$ defined in Eq. 8 and its smooth approximation $\pi_{\varepsilon}$ in Eq. 14. Assume that $\mathbf{A}^{\rm T}\mathbf{A}$ is $c$-diagonal block dominant as in 1. Suppose $\frac{\lambda}{\delta}\geq\frac{64m}{c\sqrt{\varepsilon}}$. Then, there exists a dimension-independent constant $C$ such that
@@
Here $\pi_{i}$ and $\pi_{\varepsilon,i}$ denote the marginal distributions of $\pi$ and $\pi_{\varepsilon}$ on $\mathbf{x}_{i}$, respectively, and $W_{1}$ denotes the Wasserstein-$1$ distance
$W_{1}(\mu,\nu):=\inf_{\gamma\in\Pi(\mu,\nu)}\int\left\|\mathbf{x}-\mathbf{y}\right\|_{2}\gamma(\mathbf{x},\mathbf{y})\mathrm{d}\mathbf{x}\mathrm{d}\mathbf{y},$
@@ 4 Dimension-independent block acceptance and convergence rate
It is shown in [32] that MLwG given in Algorithm 1 has dimension-independent block acceptance rate and dimension-independent convergence rate for smooth densities under assumptions on sparse conditional structure and blockwise log-concavity. In the following, we present two results which show that we can obtain similar dimension-independent results for $\pi_{\varepsilon}$ by using the concept of $c$-diagonal block dominance in 1. Since the proofs are rather similar as the ones for Proposition 3.3 and Theorem 3.6 in [32], we skip them here.
@@
Suppose $\mathbf{A}$ is bounded in the sense that
(16) $\exists C_{\mathbf{A}}>0\quad\text{s.t.}~{}\forall i,j\in[b],~{}\left\|\mathbf{A}_{ij}\right\|_{2}\leq C_{\mathbf{A}}.$
@@
Here $M$ is a dimension-independent constant depending on the block size $q$, $C_{\mathbf{A}},\delta,\varepsilon$, and $\max_{j}\left\|[\mathbf{A}\mathbf{x}^{k,i-1}-\mathbf{y}]_{j}\right\|_{2}$.
@@
Suppose $\mathbf{A}$ is bounded as in Eq. 16, $\mathbf{A}^{\rm T}\mathbf{A}$ is $c$-diagonally block dominant, and $\frac{\lambda}{\delta}\geq\frac{64m}{c\sqrt{\varepsilon}}$. Then, there exists $\tau_{0}>0$, independent of the number of blocks $b$, such that for all $0<\tau\leq\tau_{0}$, we can couple two MLwG samples $\mathbf{x}^{k}$ and $\mathbf{z}^{k}$ such that
$\sum_{j=1}^{b}\left[\mathbb{E}\left\|\mathbf{x}_{j}^{k}-\mathbf{z}_{j}^{k}\right\|_{2}\right]^{2}\leq\left(1-\lambda c\tau/4\right)^{2k}\sum_{j=1}^{b}\left[\mathbb{E}\left\|\mathbf{x}_{j}^{0}-\mathbf{z}_{j}^{0}\right\|_{2}\right]^{2}.$
Here $\tau_{0}$ ensures that $\lambda c\tau/4\in(0,1)$. In particular, one can let $\mathbf{z}^{0}\sim\pi_{\varepsilon}$, and it follows that $\mathbf{z}^{k}\sim\pi_{\varepsilon}$, which in turn shows that $\mathbf{x}^{k}$ converges to $\pi_{\varepsilon}$ geometrically fast.
@@ 5 Local & parallel MLwG algorithm
In the posterior Eq. 14, the maximal cliques can be associated to the pixels. This is due to the convolution, where all pixels within the convolution radius have a direct relationship with each other. A Gibbs density can be factorized over its maximal cliques, and thus, we can write for the posterior Eq. 14
(18) $\pi(\mathbf{x})\propto\exp{(-\sum_{\alpha\in[d]}V_{\alpha}(\mathbf{x}))},$
@@
$V_{\alpha}(\mathbf{x})=\frac{\lambda}{2}(\mathbf{y}_{\alpha}-\mathbf{A}_{\alpha}\mathbf{x})^{2}+\delta\sqrt{(\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon}.$
In fact, each clique potential $V_{\alpha}$ only depends on the pixels belonging to the clique, but for now we write $V_{\alpha}(\mathbf{x})$.
@@
Let $\Theta_{\alpha}\subseteq[d]$ be the index set containing $\alpha$ and the indices of all pixels within a frame of width $r$ around pixel $\mathbf{x}_{\alpha}$. Analogously, let $\Phi_{\alpha}\subseteq[d]$ be the index set containing $\alpha$ and the indices of all pixels within a frame of width $2r$ around pixel $\mathbf{x}_{\alpha}$. Here the integer $r>0$ is the width of the discrete PSF. Then, we can express the full conditional of $\mathbf{x}_{\alpha}$ as
(19) $\mathbf{x}_{\alpha}\sim\pi(\mathbf{x}_{\alpha}|\mathbf{x}_{\setminus\alpha})=\pi(\mathbf{x}_{\alpha}|\mathbf{x}_{\Phi_{\alpha}\setminus\alpha})\propto\exp{(-\sum_{\beta\in\Theta_{\alpha}}V_{\beta}(\mathbf{x}_{\Phi_{\alpha}}))}.$
@@
In Lemma 5.1, $\Theta_{\alpha}$ is the index set of all clique potentials that depend on $\mathbf{x}_{\alpha}$, and $\Phi_{\alpha}$ is the index set of all pixels on which the clique potentials $\{V_{\alpha}\,|\,\alpha\in\Theta_{\alpha}\}$ depend. Lemma 5.1 relies on the well-known Hammersly-Clifford theorem [11], and Eq. 19 basically states that pixel $\mathbf{x}_{\alpha}$ and all pixels in $[d]\setminus\Phi_{\alpha}$ are conditionally independent given the pixels in $\Phi_{\alpha}\setminus\alpha$. Or in other words, the full conditional of pixel $\mathbf{x}_{\alpha}$ depends only on its neighborhood given by $\Phi_{\alpha}\setminus\alpha$. These reduced conditional dependence relationships are generally known as sparse conditional structure, and we equivalently say that the posterior density Eq. 14 is local [32, 20].
Sparse conditional structure is often a prerequisite for an efficient and dimension-independent MLwG sampler. Indeed, Lemma 5.1 allows us to equivalently express the target $\pi(\mathbf{x}_{i}|\mathbf{x}_{\setminus i})$ of the blocked MLwG Algorithm 1 in Section 2.2 by a density with reduced conditional dependencies. Concretely, we can write similarly to Eq. 19
(20) $\mathbf{x}_{i}\sim\pi(\mathbf{x}_{i}|\mathbf{x}_{\setminus i})=\pi(\mathbf{x}_{i}|\mathbf{x}_{\Phi_{i}\setminus i})\propto\exp{(-\sum_{\beta\in\Theta_{i}}V_{\beta}(\mathbf{x}_{\Phi_{i}}))},$
@@
In the following sections, we outline how $\exp{(-\sum_{\beta\in\Theta_{i}}V_{\beta}(\mathbf{x}_{\Phi_{i}}))}$ and its gradient, which is required for the MALA proposal Eq. 12, can be evaluated efficiently. To this end, we define for the $(k,i-1)$-th state of the sample chain the local block likelihood and the local block prior
$\mathbf{x}_{i}^{k,i}\sim\exp\left(-l^{k,i-1}(\mathbf{x}_{i}^{k,i})\right)\quad\mathrm{and}\quad\mathbf{x}_{i}^{k,i}\sim\exp\left(-\varphi_{\varepsilon}^{k,i-1}(\mathbf{x}_{i}^{k,i})\right),$
with the potentials $l^{k,i-1}:\mathbb{R}^{q}\to\mathbb{R}_{\geq 0}$ and $\varphi_{\varepsilon}^{k,i-1}:\mathbb{R}^{q}\to\mathbb{R}_{\geq 0}$ such that
(21) $\mathbf{x}_{i}^{k,i}\sim\exp{(-\sum_{\beta\in\Theta_{i}}V_{\beta}(\mathbf{x}_{i}^{k,i},\mathbf{x}_{\Phi_{i}\setminus i}^{k,i}))}=\exp\left(-l^{k,i-1}(\mathbf{x}_{i}^{k,i})\right)\exp\left(-\varphi_{\varepsilon}^{k,i-1}(\mathbf{x}_{i}^{k,i})\right).$
In the potentials $l^{k,i-1}$ and $\varphi_{\varepsilon}^{k,i-1}$, the pixels $\mathbf{x}_{\Phi_{i}\setminus i}$ are fixed at the $(k,i-1)$-th state of the sample chain, i.e., $\mathbf{x}_{\Phi_{i}\setminus i}^{k,i}=\mathbf{x}_{\Phi_{i}\setminus i}^{k,i-1}$, such that $l^{k,i-1}$ and $\varphi_{\varepsilon}^{k,i-1}$ are functions in $\mathbf{x}_{i}^{k,i}$. In the following sections, we omit the superscripts $k,i$ of $\mathbf{x}_{i}^{k,i}$ for readability and outline separately the explicit computations of $l^{k,i-1}(\mathbf{x}_{i})$ and $\varphi_{\varepsilon}^{k,i-1}(\mathbf{x}_{i})$, and their gradients.

### v3

@@ 3 Posterior smoothing with dimension-independent error
Since MLwG requires the gradient of the log posterior density in Eqs. 11 and 12, we propose to approximate the non-smooth posterior density $\pi$ in Eq. 8 by a smooth density $\pi_{\varepsilon}$. Moreover, we show that the introduced error between $\pi$ and $\pi_{\varepsilon}$ is distributed uniformly, leading to a local dimension-independent error on the marginal distribution of any block ${\bm{x}}_{i}$.
The non-smoothness of the posterior Eq. 8 originates from the potential $\varphi_{0}({\bm{x}})={\rm TV}({\bm{x}})$ in Eq. 4. Hence, we replace $\varphi_{0}$ with a smoothed potential $\varphi_{\varepsilon}$ for some small $\varepsilon>0$, such that $\varphi_{\varepsilon}\to\varphi_{0}$ as $\varepsilon\to 0$. Various smoothing methods exist, but it is crucial to ensure that the introduced error remains small, particularly when ${\bm{x}}$ resides in a high-dimensional space. In this paper, we consider the approximation
(13) $\varphi_{\varepsilon}({\bm{x}}):=\delta\sum_{\alpha=1}^{d}\sqrt{({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}+\varepsilon},$
@@
(14) $\pi_{\varepsilon}({\bm{x}})\propto\exp\left(-\frac{\lambda}{2}\|{\bm{y}}-{\bm{A}}{\bm{x}}\|_{2}^{2}-\delta\sum_{\alpha=1}^{d}\sqrt{({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}+\varepsilon}\right).$
As we modify $\varphi_{0}$, a function in $\mathbb{R}^{d}$, the distance between $\pi$ and $\pi_{\varepsilon}$ generally depends on the dimension $d$. For instance, one can show $\text{KL}(\pi|\pi_{\varepsilon})=\mathcal{O}(d\varepsilon)$. Here, $\text{KL}(\pi|\pi_{\varepsilon})$ is the Kullback-Leibler divergence of $\pi$ from $\pi_{\varepsilon}$, see, e.g., [26]. However, when examining the marginals of $\pi$ and $\pi_{\varepsilon}$ over small blocks ${\bm{x}}_{i}$, we can show that the approximation error is independent of the full dimension and thus dimension-independent. We comment that such dimension independence is crucial for solving the image deblurring problem. It ensures that the smoothing error is evenly distributed across the image, rather than concentrating on certain pixels and creating unwanted artifacts in the image. Ensuring a uniformly distributed error is not a concept unique to image reconstruction, for other applications see [7, 35, 13].
The key to achieve the desired result lies in the sparse conditional structure. Intuitively, such local structure makes the distribution on small blocks mostly influenced by local modifications to $\varphi_{0}$, and modifications on remote blocks have little impact on it. To rigorously justify the above observations, we introduce the concept of $c$-diagonal block dominance, which is imposed on ${\bm{C}}={\bm{A}}^{\rm T}{\bm{A}}$ to quantify how locally concentrated the blurring matrix ${\bm{A}}$ is.
@@
Consider a positive definite matrix ${\bm{C}}\in\mathbb{R}^{d\times d}$ and let ${\bm{C}}_{ij}$ denote the $(i,j)$-th sub block of ${\bm{C}}$ of size $q$. ${\bm{C}}$ is called $c$-diagonal block dominant for some $c>0$, if there exists a symmetric matrix $M\in\mathbb{R}^{b\times b}$ s.t.
@@
[34] introduces a similar blockwise log-concavity condition, which is crucial for the dimension-independent convergence rate therein. The $c$-diagonal block dominance here can be viewed as an $\ell_{1}$ version of the blockwise log-concavity condition (which is an $\ell_{2}$ condition).
Now we state the main theorem, which shows that the marginal error of $\pi$ and $\pi_{\varepsilon}$ is dimension-independent if ${\bm{A}}^{\rm T}{\bm{A}}$ is $c$-diagonal block dominant.
@@
Consider the target distribution $\pi$ defined in Eq. 8 and its smooth approximation $\pi_{\varepsilon}$ in Eq. 14. Assume that ${\bm{A}}^{\rm T}{\bm{A}}$ is $c$-diagonal block dominant as in 1. Suppose $\frac{\lambda}{\delta}\geq\frac{64m}{c\sqrt{\varepsilon}}$. Then, there exists a dimension-independent constant $C$ such that
@@
Here, $\pi_{i}$ and $\pi_{\varepsilon,i}$ denote the marginal distributions of $\pi$ and $\pi_{\varepsilon}$ on ${\bm{x}}_{i}$, respectively, and $W_{1}$ denotes the Wasserstein-$1$ distance
$W_{1}(\mu,\nu):=\inf_{\gamma\in\Pi(\mu,\nu)}\int\left\|{\bm{x}}-{\bm{y}}\right\|_{2}\gamma({\bm{x}},{\bm{y}})\mathrm{d}{\bm{x}}\mathrm{d}{\bm{y}},$
@@ 4 Dimension-independent block acceptance and convergence rate
It is shown in [34] that MLwG given in Algorithm 1 has dimension-independent block acceptance rates and a dimension-independent convergence rate for smooth densities under assumptions on sparse conditional structure and blockwise log-concavity. In the following, we present two results, which show that we can obtain similar dimension-independent results for $\pi_{\varepsilon}$ by using the concept of $c$-diagonal block dominance in 1. Since the proofs are similar to those for Proposition 3.3 and Theorem 3.6 in [34], we skip them here.
@@
Suppose ${\bm{A}}$ is bounded in the sense that
(16) $\exists C_{\bm{A}}>0\quad\text{s.t.}~{}\forall i,j\in[b],~{}\left\|{\bm{A}}_{ij}\right\|_{2}\leq C_{\bm{A}}.$
@@
Here, $M$ is a dimension-independent constant depending on the block size $q$, $C_{\bm{A}},\delta,\varepsilon$, and $\max_{j}\left\|[{\bm{A}}{\bm{x}}^{k,i-1}-{\bm{y}}]_{j}\right\|_{2}$.
Remark 2.
In fact, the lower bound depends on the state of x through the term $\max_{j}\left\|[{\bm{A}}{\bm{x}}^{k,i-1}-{\bm{y}}]_{j}\right\|_{2}$. However, since the pixels values are bounded in practice, this term is usually also bounded and dimension-independent.
@@
Suppose ${\bm{A}}$ is bounded as in Eq. 16, ${\bm{A}}^{\rm T}{\bm{A}}$ is $c$-diagonally block dominant, and $\frac{\lambda}{\delta}\geq\frac{64m}{c\sqrt{\varepsilon}}$. Then, there exists $\tau_{0}>0$, independent of the number of blocks $b$, such that for all $0<\tau\leq\tau_{0}$, we can couple two MLwG samples ${\bm{x}}^{k}$ and ${\bm{z}}^{k}$ such that
$\sum_{j=1}^{b}\left[\mathbb{E}\left\|{\bm{x}}_{j}^{k}-{\bm{z}}_{j}^{k}\right\|_{2}\right]^{2}\leq\left(1-\lambda c\tau/4\right)^{2k}\sum_{j=1}^{b}\left[\mathbb{E}\left\|{\bm{x}}_{j}^{0}-{\bm{z}}_{j}^{0}\right\|_{2}\right]^{2}.$
Here, $\tau_{0}$ ensures that $\lambda c\tau/4\in(0,1)$. In particular, one can let ${\bm{z}}^{0}\sim\pi_{\varepsilon}$, and it follows that ${\bm{z}}^{k}\sim\pi_{\varepsilon}$, which in turn shows that ${\bm{x}}^{k}$ converges to $\pi_{\varepsilon}$ geometrically fast.
Remark 3.
The convergence result is based on the maximal coupling of two MLwG chains as in [34]. In brief, the two chains are coupled to share in each step the same $\xi_{i}^{k,i}$ in (12) and the random variable $u\sim\text{Uniform}(0,1)$ used to determine the acceptance of proposals (See Algorithm 1).
@@ 5 Local & parallel MLwG algorithm
In the posterior Eq. 14, the maximal cliques can be associated to the pixels. This is due to the convolution, where all pixels within the convolution radius have a direct relationship with each other. A Gibbs density can be factorized over its maximal cliques, and thus we can write for the posterior Eq. 14
(18) $\pi({\bm{x}})\propto\exp{\Big{(}-\sum_{\alpha\in[d]}V_{\alpha}({\bm{x}})\Big{)}},$
@@
$V_{\alpha}({\bm{x}})=\frac{\lambda}{2}({\bm{y}}_{\alpha}-{\bm{A}}_{\alpha}{\bm{x}})^{2}+\delta\sqrt{({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}+\varepsilon}.$
In fact, each clique potential $V_{\alpha}$ only depends on the pixels belonging to the clique, but for now we write $V_{\alpha}({\bm{x}})$.
@@
Let $\Theta_{\alpha}\subseteq[d]$ be the index set containing $\alpha$ and the indices of all pixels within a frame of radius $r$ around pixel ${\bm{x}}_{\alpha}$. Analogously, let $\Phi_{\alpha}\subseteq[d]$ be the index set containing $\alpha$ and the indices of all pixels within a frame of radius $2r$ around pixel ${\bm{x}}_{\alpha}$. Here, the integer $r>0$ is the radius of the discrete PSF. Then, we can express the full conditional of ${\bm{x}}_{\alpha}$ as
(19) ${\bm{x}}_{\alpha}\sim\pi({\bm{x}}_{\alpha}|{\bm{x}}_{\setminus\alpha})=\pi({\bm{x}}_{\alpha}|{\bm{x}}_{\Phi_{\alpha}\setminus\alpha})\propto\exp{\Big{(}-\sum_{\beta\in\Theta_{\alpha}}V_{\beta}({\bm{x}}_{\Phi_{\alpha}})\Big{)}}.$
@@
In Lemma 5.1, $\Theta_{\alpha}$ is the index set of all clique potentials that depend on ${\bm{x}}_{\alpha}$, and $\Phi_{\alpha}$ is the index set of all pixels on which the clique potentials $\{V_{\alpha}\,|\,\alpha\in\Theta_{\alpha}\}$ depend. Lemma 5.1 relies on the well-known Hammersly-Clifford theorem [12], and Eq. 19 basically states that pixel ${\bm{x}}_{\alpha}$ and all pixels in $[d]\setminus\Phi_{\alpha}$ are conditionally independent given the pixels in $\Phi_{\alpha}\setminus\alpha$. Or in other words, the full conditional of pixel ${\bm{x}}_{\alpha}$ depends only on its neighborhood given by $\Phi_{\alpha}\setminus\alpha$. These reduced conditional dependence relationships are generally known as sparse conditional structure, and we equivalently say that the posterior density Eq. 14 is local [21, 34].
Sparse conditional structure is often a prerequisite for an efficient and dimension-independent MLwG sampler. Indeed, Lemma 5.1 allows us to equivalently express the target $\pi({\bm{x}}_{i}|{\bm{x}}_{\setminus i})$ of the blocked MLwG Algorithm 1 by a density with reduced conditional dependencies. Concretely, we can write similarly to Eq. 19
(20) ${\bm{x}}_{i}\sim\pi({\bm{x}}_{i}|{\bm{x}}_{\setminus i})=\pi({\bm{x}}_{i}|{\bm{x}}_{\Phi_{i}\setminus i})\propto\exp{\Big{(}-\sum_{\beta\in\Theta_{i}}V_{\beta}({\bm{x}}_{\Phi_{i}})\Big{)}},$
@@
In the following sections, we outline how $\exp{(-\sum_{\beta\in\Theta_{i}}V_{\beta}({\bm{x}}_{\Phi_{i}}))}$ and its gradient, which is required for the MALA proposal Eq. 12, can be evaluated efficiently. To this end, for the $(k,i-1)$-th state of the sample chain, we define
  • the local block likelihood: $\exp\left(-l^{k,i-1}({\bm{x}}_{i}^{k,i})\right)$ and
  • the local block prior: $\exp\left(-\varphi_{\varepsilon}^{k,i-1}({\bm{x}}_{i}^{k,i})\right)$
with the potentials $l^{k,i-1}:\mathbb{R}^{q}\to\mathbb{R}_{\geq 0}$ and $\varphi_{\varepsilon}^{k,i-1}:\mathbb{R}^{q}\to\mathbb{R}_{\geq 0}$, such that
(21) ${\bm{x}}_{i}^{k,i}\sim\exp{\Big{(}-\sum_{\beta\in\Theta_{i}}V_{\beta}({\bm{x}}_{i}^{k,i},{\bm{x}}_{\Phi_{i}\setminus i}^{k,i})\Big{)}}=\exp\left(-l^{k,i-1}({\bm{x}}_{i}^{k,i})-\varphi_{\varepsilon}^{k,i-1}({\bm{x}}_{i}^{k,i})\right).$
In the potentials $l^{k,i-1}$ and $\varphi_{\varepsilon}^{k,i-1}$, the pixels ${\bm{x}}_{\Phi_{i}\setminus i}$ are fixed at the $(k,i-1)$-th state of the sample chain, i.e., ${\bm{x}}_{\Phi_{i}\setminus i}^{k,i}={\bm{x}}_{\Phi_{i}\setminus i}^{k,i-1}$, such that $l^{k,i-1}$ and $\varphi_{\varepsilon}^{k,i-1}$ are functions in ${\bm{x}}_{i}^{k,i}$. In the following sections, we omit the superscripts $k,i$ of ${\bm{x}}_{i}^{k,i}$ for readability and outline separately the explicit computations of $l^{k,i-1}({\bm{x}}_{i})$ and $\varphi_{\varepsilon}^{k,i-1}({\bm{x}}_{i})$, and their gradients.